Risk-free investment pays $550 in two years, risk-free interest rate is 5%. What is the highest...

  1. Risk-free investment pays $550 in two years, risk-free interest rate is 5%. What is the highest amount you would agree to pay for this security?

  1. Security’s market price is $500, risk-free rate is 5%. The security makes one risk-free payment in one year and matures. What must be risk-free payoff if the security in one year to make you interested in investing in this security?
Answer #1

IN FIRST CASE, WE HAVE TO FIND OUT, PRESENT VALUE :

PV = FV/(1+r)n

PV = 550/(1+0.05)2 = 550/1.1025 = 498.8662 = 498.87

ANSWER : HIGHEST AMOUNT AGREE TO PAY = 498.87

IN SECOND CASE, WE HAVE TO FIND OUT, FUTURE VALUE :

FV = PV*(1+r)n

PV = 500*(1+0.05) = 525

ANSWER : HIGHEST AMOUNT AGREE TO PAY = 525 (THUMBS UP PLEASE)